OpenStudy (midhun.madhu1987):

Prove the following by using the principle of mathematical induction: x^(2n) - y^(2n) is divisible by (x+y) I tried to do, but i'm not getting. So need the steps.

OpenStudy (mokeira):

when n=1, the given statement becomes \[x ^{2}-y ^{2} \] is divisible by (x+y) therefore the statement is true

OpenStudy (mokeira):

I can give you an article on the same if you want

OpenStudy (midhun.madhu1987):

it should be true for all values of n. Not 1 alone...

OpenStudy (mokeira):

choose a value and let me prove it

OpenStudy (midhun.madhu1987):

I need to prove it using the Principle of Mathematical Induction. First prove it for n=1. Then assume it for n=k. then prove n=k+1 is true for every n=k.

OpenStudy (mokeira):

ok

OpenStudy (mokeira):

I see what you mean

OpenStudy (midhun.madhu1987):

Any help would be much appreciated.. as it took my most of the time.. still didn't find a comfortable way to prove it.. :(

OpenStudy (mokeira):

is it x+y or x-y?

OpenStudy (aum):

Assume it is true for k. That is, \(x^{2k} - y^{2k}\) is divisible by (x+y) ----- (1) Then, for k + 1, the numerator is: \(x^{2(k+1)} - y^{2(k+1)} = x^2x^{2k} - y^2y^{2k} = x^2x^{2k} - (y^2+x^2-x^2)y^{2k} = \\ x^2x^{2k} - x^2y^{2k} - (y^2-x^2)y^{2k} = x^2(x^{2k}-y^{2k}) - (y+x)(y-x)y^{2k} \) . The first term is divisible by (x+y) due to (1) and the second term is divisible by (x+y) due to the factor (x+y). Therefore, if it is true for k, it is true for (k+1).

OpenStudy (midhun.madhu1987):

x+y

OpenStudy (midhun.madhu1987):

@aum Great Job.. Thanks... is this the only method??

OpenStudy (aum):

For proof by induction this is what occurred to me. Others may have a different proof but I can't imagine t would be any simpler than this one as far as induction proof is concerned.

ganeshie8 (ganeshie8):

cleverly done xD

OpenStudy (anonymous):

@ganeshie8 plz zhelp me sir..i have messged u